Selected transactions completed by Primo Discount Corporation during the current fiscal year are as follows:

Jan. 9 Split the common stock 3 for 1 and reduced the par from $75 to $25 per share. After the split, there were 1,279,500 common shares outstanding.
Feb. 28 Purchased 39,200 shares of the corporation’s own common stock at $27, recording the stock at cost.
May 1 Declared semiannual dividends of $0.90 on 75,000 shares of preferred stock and $0.15 on the common stock to stockholders of record on June 1, payable on July 10.
Jul. 10 Paid the cash dividends.
Sep. 7 Sold 30,800 shares of treasury stock at $34, receiving cash.
Oct. 1 Declared semiannual dividends of $0.90 on the preferred stock and $0.15 on the common stock (before the stock dividend). In addition, a 3% common stock dividend was declared on the common stock outstanding. The fair market value of the common stock is estimated at $38.
Dec. 1 Paid the cash dividends and issued the certificates for the common stock dividend.

Journalize the transactions. If no entry is required, simply skip to the next transaction. Refer to the Chart of Accounts for exact wording of account titles.
Answer #1

Journal Entries:

Date Account Titles and Explanations Debit Credit
Jan. 9 No Journal Entry is required for share split because there is no change in the common stock due to stock split as the total value remains same.
Feb. 28 Treasury Stock (39,200 shares * $27 per share) $1,058,400
   Cash $1,058,400
(To record the purchase of treasury stock)
May. 1 Preferred Dividends (75,000*$0.90) $67,500
Common Dividends (1,279,500 - 39,200 = 1,240,300 * $0.15) $186,045
   Dividends Payable $253,545
(To record the declaration of preferred and common dividends)
Jul. 10 Dividends Payable $253,545
   Cash $253,545
(To record the payment of dividends)
Sep. 1 Cash (30,800 shares * $34 per share) $1,047,200
   Common Stock (30,800 shares * $25 per share) $770,000
   Paid-in Capital from Sale of Treasury Stock ($34 - $25 = $9 * 30,800) $277,200
(To record the sale of treasury stock)
Oct. 1 Preferred Dividends (75,000*$0.90) $52,800
Common Dividends (1,279,500 - 39,200 + 30,800 = 1,271,100 * $0.15) $190,665
   Dividends Payable $243,465
(To record the declaration of preferred and common dividends)
Oct. 1 Stock Dividends (1,271,100*3/100 = 38,133 shares * $38 per share) $1,449,054
   Common Stock Dividends Distributable $1,449,054
(To record the declaration of 3% stock dividends)
Dec. 1 Dividends Payable $243,465
   Cash $243,465
(To record the payment of dividends)
Dec. 1 Common Stock Dividends Distributable $1,449,054
   Common Stock (38,133 shares * $25 per share) $953,325
   Paid-in Capital in Excess of Par-Common Stock ($38 - $25 = $13 * 38,133) $495,729
(To record the payment of stock dividends)